OpenStudy (anonymous):

Math test questions! :) Only 5 of them, I'll take screenshots and give you the pictures.

OpenStudy (abb0t):

NO CHEATING!!!!

OpenStudy (anonymous):

I already put down an answer to them, if I wanted to cheat I'd review the test answers and take it with them. I need conformation that they're correct, that's all :)

OpenStudy (anonymous):

ill try and help!

OpenStudy (anonymous):

Thanks! let me show you the first question haha

OpenStudy (anonymous):

Oh, I just realized you see what I named the file, LOL

OpenStudy (anonymous):

is 78 correct? I'm pretty sure I don't have many right answers at all in this quiz.

sammixboo (sammixboo):

Well, does \(78~=\frac{78}{13}\)?

OpenStudy (anonymous):

No?

sammixboo (sammixboo):

It doesn't. To solve this you must do \(\frac{78}{13}\)

OpenStudy (anonymous):

I can't do fractions, so I'm gonna fail at this, sorry

sammixboo (sammixboo):

Well just do 78 divided by 13

OpenStudy (anonymous):

that's 6

sammixboo (sammixboo):

Yes, so your height is 6

OpenStudy (anonymous):

yes somtimes its not a fraction lol

OpenStudy (anonymous):

So the final answer for this one is 6?

OpenStudy (anonymous):

mhm!

sammixboo (sammixboo):

Yes. \(\frac{78}{13}\) = 78 divided by 13

OpenStudy (anonymous):

Thank you guys! :) I have 4 more, would you have the time for that?

sammixboo (sammixboo):

I will help you check your answers :)

OpenStudy (anonymous):

same here!

OpenStudy (anonymous):

Thank you guys so much in advance :)

sammixboo (sammixboo):

Does \(6.8~=\frac{2\times17}{5}\) (also know as) \(6.8~=\frac{34}{5}\)

sammixboo (sammixboo):

Does 34 divided by 5 equal 6.8?

OpenStudy (anonymous):

Yes, it does

sammixboo (sammixboo):

Yes! Corrct, so that is your answer

OpenStudy (anonymous):

Thanks! Alright, question 3 :D

sammixboo (sammixboo):

Does \(435.2~=\frac{25.6}{8}\)

OpenStudy (anonymous):

I'm not sure, probably not?

sammixboo (sammixboo):

Nope :( All you need to so is divide 25.6 by 8, so what is 25.6/8?

OpenStudy (anonymous):

3.2!

sammixboo (sammixboo):

Correct!

OpenStudy (anonymous):

Almost done, omg

OpenStudy (anonymous):

your bettingme to it! ^_^

sammixboo (sammixboo):

Does \(25~=~26.5\times1.5\)?

OpenStudy (anonymous):

Not at all

OpenStudy (anonymous):

Unless 25 magically became 39.75?

OpenStudy (anonymous):

Which is highly unlikely.

sammixboo (sammixboo):

Nope! Here is our formula \(d~=~rt\) Plug in the info given. \(25~=~r\times 1.5\) To solve this we must divide 1.5 by 25, so what is 25/1.5?

sammixboo (sammixboo):

Round to the nearest tenth

OpenStudy (anonymous):

16.67 or 16.6

sammixboo (sammixboo):

Or 16.7

OpenStudy (anonymous):

going with 16.67

OpenStudy (anonymous):

Or 16/7, yeah, lol

sammixboo (sammixboo):

:) Good job!

sammixboo (sammixboo):

Ahaha typos

OpenStudy (anonymous):

Alright, last question, you made it seem really easy by breaking it up like that, lol :)

sammixboo (sammixboo):

Haha :)

sammixboo (sammixboo):

\(P~=~ns\) \(54.4~=~8s\) To solve this you must divide 8 on both sides of the equation. \(54.4/8~=~8s/8\) \(54.4/8~=~s\) So what is 54.4/8?

OpenStudy (anonymous):

6.8! :)

sammixboo (sammixboo):

Correct! Good job :) If you need any more help tag me @sammixboo

OpenStudy (anonymous):

Thank you so much! I got a 100. Night, :)